OpenStudy (saucvi):

Can someone check my work for this? Prove that a line that divides two sides of a triangle proportionally is parallel to the third side. Be sure to create and name the appropriate geometric figures.

OpenStudy (saucvi):

My answer: Triangle DEF has two midpoints. Midpoint G between points D and F. Midpoint H between points D and E. Segment GH is parallel to segment FE. Using the triangle proportionality theorem you can determine the triangles corresponding. The triangle proportionality theorem is AB/BC = AE/DE. If HE = 20, DG = 24, and GF = 12 you can find the measurement of DH with the triangle proportionality theorem. DH/20 = 24/12 DH/20 = 2 DH = 40.

OpenStudy (saucvi):



OpenStudy (bobo-i-bo):

Yes, that's fine

OpenStudy (saucvi):

Alright, thank you!